OpenStudy (darkbluechocobo):

Help check work with Writing a Vector as the Sum of Two Vector Components

OpenStudy (darkbluechocobo):

Write v as the sum of two vector components if v = -5i + 6j and w = 3i + 4j

OpenStudy (darkbluechocobo):

(-5*3)+(6*4) -15+24 9

OpenStudy (darkbluechocobo):

\[\frac{ 9 }{ \sqrt(6)^2+(4)^2 }=\frac{ 9 }{ 25 }(3,4)\]

OpenStudy (darkbluechocobo):

v1= (27/25,36/25)

OpenStudy (darkbluechocobo):

v2=v-v1= -5i + 6 - (27/25,36/25)

OpenStudy (darkbluechocobo):

v2=(-152/25, 114,25) v= (27/25i + 36/25j)+(-152/25i +114,25j)

OpenStudy (anonymous):

too many problems, why did you multiply in the first

OpenStudy (darkbluechocobo):

hm?

OpenStudy (darkbluechocobo):

What do you mean

OpenStudy (darkbluechocobo):

@zepdrix

zepdrix (zepdrix):

hmm i don't understand the question 0_o

zepdrix (zepdrix):

as the sum of two unit vectors or something? what does this have to do with w?

OpenStudy (darkbluechocobo):

Well to Write v as the sum of two vector components. V1= the projection from v to w. V2 = v-v1

OpenStudy (darkbluechocobo):

Thats why the w is needed :p

zepdrix (zepdrix):

vector projection? :o mmm lil rusty.. thinking

zepdrix (zepdrix):

bah i dunno >.<